how do you get the y intercept form for

-x +2y = -2
this comes from the nort carolina algebra 1 prentice hall mathmatics book just so u know I need help by 9:00pm today please help me

Answer by checkley71(8403)

-X+2Y=-2 OR 2Y=X-2 OR DIVIDING BOTH SIDES BY 2 WE GET Y=X/2-1 THEREFORE -1 IS THE Y INTERCEPT